AnimeDude743:

0.01x - 0.3y = 1 y = 0.1x - 2 Which of the following systems is equivalent to the given system? x - 30y = 100 and 10y = x - 20 10x - 30y = 10 and 10y = x - 20 x - 3y = 10 and 10y = x - 20

SmokeyBrown:

If you multiply every element in an equation by the same constant, the result is "equivalent" to the original, since the relationship between the elements is preserved. For example, you can multiply every element in the first equation 0.01x - 0.3y = 1 by 100 to get x - 30y = 100 Similarly, you can multiply every element in the second equation y = 0.1x - 2 by 10 to get 10y = x - 20 So, the first option would be an equivalent set of equations to the original.